OpenStudy (anonymous):

How do you solve (g+4)-3G=1+g

OpenStudy (anonymous):

you take out the parenthesis around g+4 because there's nothing to be distributed to it. g+4-3g=1+g -4 -4 g-3g=-3+g -g -g -3g=-3 /-3 /-3 g=1 check your work and pluck in 1 instead of g in the original equation.